i m thinking of 3 consecutive odd integer. the sum of the first and last of these integers is 26. what are the integers
Answer by elima(1433)

i m thinking of 3 consecutive odd integer. the sum of the first and last of these integers is 26. what are the integers
let x be the first integer;
x+1; second integer
x+2 third integer
x+x+2=26
2x+2=26
2x=24
x=12; first integer
12+1=13; second integer
12+2=14; third integer
:)